# Weighted shift operators on spaces of analytic functions over an annulus - I

### v1

@@ 3.2. Linear dynamics of $F_{w}$ and its adjoint $F_{w}^{*}$
    $\sum_{n=1}^{\infty}\left\lvert\frac{a_{n}}{w_{0}w_{1}\cdots w_{n-1}}\right\rvert^{p}<\infty$ and $\sum_{n=1}^{\infty}\lvert w_{-1}w_{-2}\cdots w_{-n}b_{-n-1}\rvert^{p}<\infty.$
Then $(iii)\Rightarrow(i)\Rightarrow(ii).$
@@
(i) implies (ii) trivial.
To see that (iii) implies (i), suppose that $\sum_{n=1}^{\infty}\left\lvert\frac{a_{n}}{w_{0}\cdots w_{n-1}}\right\rvert^{p}<\infty$ and $\sum_{n=1}^{\infty}\lvert w_{-1}\cdots w_{-n}b_{-n-1}\rvert^{p}<\infty$ hold. We apply the chaoticity criterion to show that $F_{w}^{*}$ is chaotic on $(\ell^{p}_{a,b}(\mathtt{A}))^{*}.$ Consider $D=\text{span}\{k_{n}(z):n\in\mathbb{Z}\}$. Then by the Propostion 2.3, the set $D$ is dense in $(\ell^{p}_{a,b}(\mathtt{A}))^{*}.$ Define $S:D\rightarrow D$ by $S(k_{n})=\frac{1}{w_{n}}k_{n+1},~{}~{}\forall~{}~{}n\in\mathbb{Z}.$ Since $F_{w}^{*}S(f)=f,$ $\forall$ $f\in D,$ we just need to show that $\sum_{n=0}^{\infty}(F_{w}^{*})^{n}(f)$ and $\sum_{n=0}^{\infty}S^{n}(f)$ converge unconditionally, for each $f\in D.$ We prove that $\sum_{n=1}^{\infty}w_{-1}w_{-2}\cdots w_{-n}k_{-n}$ and $\sum_{n=1}^{\infty}\frac{1}{w_{0}w_{1}\cdots w_{n-1}}k_{n}$ are unconditionally convergent in $(\ell^{p}_{a,b}(\mathtt{A}))^{*}.$ Since $k_{n}=a_{n}f_{n}^{*}+b_{n-1}f_{n-1}^{*}$, for all $n\in\mathbb{Z}$, we have
@@
for all $n\geq N$. The right hand side of the above inequality is the $\ell^{p}$ element $\left\{\frac{a_{n+1}}{w_{0}\cdots w_{n}}\right\}_{n=1}^{\infty},$ and hence it is absolutely $p$-summable. Consequently, the series $\sum_{n\geq N}\frac{1}{w_{0}w_{1}\cdots w_{n-1}}k_{n}$ is convergent. The unconditional convergence follows because $\{f_{n}^{*}\}$ is an unconditional basis. Similarly, to show that $\sum_{n=1}^{\infty}w_{-1}w_{-2}\cdots w_{-n}k_{-n}$ is unconditionally convergent, we write
@@
As $\limsup_{n}|a_{-n}/b_{-n-1}|<1$, one gets $N\in\mathbb{N}$ and $R<1$ such that $|a_{-n}/b_{-n-1}|<R$ for all $n\geq N$. Hence,
@@
for all $n\geq N$. The right hand side of the above inequality is the $\ell^{p}$ element $\left\{w_{-1}\cdots w_{-n}b_{-n-1}\right\}_{n\geq 1}$ and hence it is absolutely $p$-summable. Consequently, the series $\sum_{n\geq N}\frac{1}{w_{-1}w_{-2}\cdots w_{-n}}k_{-n}$ is unconditionally convergent. Hence, $F_{w}^{*}$ satisfies the chaoticity criterion, and (i) follows.
@@ 3.3. Examples and Illustrations
We illustrate the above results on dynamics with explicit examples. Recall that a Laurent series $\sum_{n=-\infty}^{\infty}a_{n}z^{n}$ converges for all $z$ such that $r<\lvert z\rvert<R,$ where
@@
Let $a_{0}=b_{0}=w_{0}=1,$ and $a_{n}=1,$ $b_{n}=\frac{1}{n},$ $w_{n}=\frac{1}{(n+1)!},$ for all $n\geq 1,$, along with $a_{n}=\frac{n2^{n}}{2^{2}},$ $b_{n}=n2^{n},$ $w_{n}=8,$ for all $n\leq-1.$ Now by Theorem 2.1, for $1<p<\infty,$
$\sum_{n=-\infty}^{\infty}\left((\lvert a_{n}\rvert+\lvert b_{n}\rvert)\lvert z\rvert^{n}\right)^{q}=\sum_{n=1}^{\infty}\left(\left(\frac{n}{2^{n}}\frac{1}{2^{2}}+\frac{n}{2^{n}}\right)\lvert z\rvert^{-n}\right)^{q}+\sum_{n=0}^{\infty}\left(\left(1+\frac{1}{n}\right)\lvert z\rvert^{n}\right)^{q},$
where $11/p+1/q=$ 1. A simple computation shows that the natural annulus $\mathtt{A}$ of $\ell^{p}_{a,b}(\mathtt{A})$ is $r<|z|<R$, where $r=\frac{1}{2}$ and $R=1.$ Therefore $\sum_{n=-\infty}^{\infty}\left((\lvert a_{n}\rvert+\lvert b_{n}\rvert)\lvert z\rvert^{n}\right)^{q}<\infty$ if $\frac{1}{2}<\lvert z\rvert<1.$ In view of Theorem 3.4, $F_{w}$ is bounded on $\ell^{p}_{a,b}(\mathtt{A}),$ where $\mathtt{A}$ is the given by annulus $\frac{1}{2}<\lvert z\rvert<1.$ Additionally, $F_{w}$ satisfies the conditions of Theorems 3.7 and 3.9, and so $F_{w}$ is mixing and chaotic on $\ell^{p}_{a,b}(\mathtt{A}).$
@@
In the previous example, if we let $w_{n}=4,$ for all $n\leq-1,$ $w_{n}=\frac{1}{(n+1)!},$ for all $n\geq 0,$ and $\{a_{n}\},\{b_{n}\}$ being the same as mentioned above, then we see that by Theorem 3.7, $F_{w}$ is supercyclic but not hypercyclic on $\ell^{p}_{a,b}(\mathtt{A})$ as
$\lim_{n\rightarrow\infty}\left\lvert\frac{(w_{\nu-1}\cdots w_{\nu-n})(a_{-1}\cdots a_{\nu-n})}{b_{-1}\cdots b_{\nu-n}}\right\rvert<\infty,$
for $\nu\in\mathbb{N}$ with $n>\nu.$
@@
$\sum_{n=-\infty}^{\infty}\left((\lvert a_{n}\rvert+\lvert b_{n}\rvert)\lvert z\rvert^{n}\right)^{q}=\sum_{n=1}^{\infty}\left(\left(\frac{1}{n}+1\right)\lvert z\rvert^{-n}\right)^{q}+\sum_{n=0}^{\infty}\left(\left(\frac{n}{2^{n}}+\frac{n}{2^{n+2}}\right)\lvert z\rvert^{n}\right)^{q},$
where $11/p+1/q=$ 1. In this case, $r=1$, $R=2$, and hence, the annulus $\mathtt{A}$ is $21<|z|<$ 2. Further, $\sum_{n=-\infty}^{\infty}\left((\lvert a_{n}\rvert+\lvert b_{n}\rvert)\lvert z\rvert^{n}\right)^{q}<\infty$ if $21<\lvert z\rvert<$ 2 . In addition, by the conditions in Theorem 3.4, the weighted shift $F_{w}$ is bounded on $\ell^{p}_{a,b}(\mathtt{A}).$ In view of Theorems 3.10 and 3.11, the adjoint $F_{w}^{*}$ is mixing and chaotic on $\ell^{p}_{a,b}(\mathtt{A})^{*}.$ However, the $F_{w}$ itself is not supercyclic.

### v2

@@ 3.2. Linear dynamics of $F_{w}$ and its adjoint $F_{w}^{*}$
    $\sum_{n=1}^{\infty}\left\lvert\frac{a_{n}}{w_{0}w_{1}\cdots w_{n-1}}\right\rvert^{q}<\infty$ and $\sum_{n=1}^{\infty}\lvert w_{-1}w_{-2}\cdots w_{-n}b_{-n-1}\rvert^{q}<\infty.$
Then $(iii)\Rightarrow(i)\Rightarrow(ii).$ In addition, if for $\frac{1}{p}+\frac{1}{q}=1,$
$\sum_{n\geq 0}\left(\sum_{j\geq 1}\left\lvert\frac{\prod_{k=1}^{j}b_{n+k-1}}{\prod_{k=1}^{j}a_{n+k}}\right\rvert^{p}\right)^{\frac{q}{p}}<\infty~{}~{}~{}~{}~{}~{}~{}~{}\text{and}~{}~{}~{}~{}~{}~{}~{}~{}\sum_{n\geq 0}\left(\sum_{j\geq 1}\left\lvert\frac{\prod_{k=1}^{j}a_{-n-k+1}}{\prod_{k=1}^{j}b_{-n-k}}\right\rvert^{p}\right)^{\frac{q}{p}}<\infty,$
then $(ii)\Rightarrow(iii)$.
@@
It is trivial that (i) implies (ii).
To see that (ii) implies (iii), let $\psi=\sum_{m=-\infty}^{\infty}\lambda_{m}f_{m}^{*}$ be a non-zero periodic vector for $F_{w}^{*}$ on $(\ell^{p}_{a,b}(\mathtt{A}))^{*}$, where $f_{m}^{*}$, $m\in\mathbb{Z}$, forms an unconditional Schauder basis for $(\ell^{p}_{a,b}(\mathtt{A}))^{*}$. Now by using Proposition 3.11, we get
$\psi=\sum_{m=-\infty}^{\infty}\lambda_{m}f_{m}^{*}:=\sum_{m=-\infty}^{\infty}A_{m}k_{m},$
where
$A_{0}=\displaystyle\sum_{j=2}^{\infty}(-1)^{j-1}\lambda_{-j}\frac{a_{-1}a_{-2}\cdots a_{-j+1}}{b_{-1}b_{-2}\cdots b_{-j}}+\displaystyle\sum_{j=1}^{\infty}(-1)^{j}\lambda_{j}\frac{b_{0}b_{1}\cdots b_{j-1}}{a_{0}a_{1}\cdots a_{j}},$
$A_{m}=\lambda_{m}\frac{1}{a_{m}}+\displaystyle\sum_{j=1}^{\infty}(-1)^{j}\lambda_{m+j}\frac{b_{m}b_{m+1}\cdots b_{m+j-1}}{a_{m}a_{m+1}\cdots a_{m+j}},$
and
$A_{-m}=\lambda_{-m-1}\frac{1}{b_{-m-1}}+\displaystyle\sum_{j=1}^{\infty}(-1)^{j}\lambda_{-m-1-j}\frac{a_{-m-1}a_{-m-2}\cdots a_{-m-j}}{b_{-m-1}b_{-m-2}\cdots b_{-m-1-j}},$
for all $m\in\mathbb{N}.$ Now, let $\nu\in\mathbb{N}$ such that $(F_{w}^{*})^{\nu}\psi=\psi$ for $z\in\mathtt{A}$. Then,
$(F_{w}^{*})^{n\nu}\psi=\psi$ for all $n\geq 1$.
Equating the coefficient of $k_{0}$ both sides, we get
$\frac{w_{n\nu-1}\cdots w_{0}}{a_{n\nu}}\left(\lambda_{n\nu}+\sum_{j=1}^{\infty}(-1)^{j}\lambda_{n\nu+j}\frac{b_{n\nu}\cdots b_{n\nu+j-1}}{a_{n\nu+1}\cdots a_{n\nu+j}}\right)=\cdots-\lambda_{-2}\frac{a_{-1}}{b_{-2}b_{-1}}-\lambda_{1}\frac{b_{0}}{a_{0}a_{1}}+\lambda_{2}\frac{b_{0}b_{1}}{a_{0}a_{1}a_{2}}-\cdots,$
Since the series in the right hand side converges, we get that
$C\sum_{n=1}^{\infty}\left\lvert\frac{a_{n\nu}}{w_{n\nu-1}w_{n\nu-2}\cdots w_{0}}\right\rvert^{q}=\sum_{n=1}^{\infty}\left\lvert\lambda_{n\nu}+\sum_{j=1}^{\infty}(-1)^{j}\lambda_{n\nu+j}\frac{b_{n\nu}\cdots b_{n\nu+j-1}}{a_{n\nu+1}\cdots a_{n\nu+j}}\right\rvert^{q},$
for some $C>0$. Note that using Hölder inequality, we can get that
$\sum_{n=1}^{\infty}\left\lvert\lambda_{n\nu}+\sum_{j=1}^{\infty}(-1)^{j}\lambda_{n\nu+j}\frac{b_{n\nu}\cdots b_{n\nu+j-1}}{a_{n\nu+1}\cdots a_{n\nu+j}}\right\rvert^{q}<\infty,$
as $\{\lambda_{n}\}_{n}\in\ell^{q}(\mathbb{Z})$ and $\sum_{n\geq 0}\left(\sum_{j\geq 1}\left\lvert\frac{\prod_{k=1}^{j}b_{n+k-1}}{\prod_{k=1}^{j}a_{n+k}}\right\rvert^{p}\right)^{\frac{q}{p}}<\infty.$ Therefore,
$\sum_{n=1}^{\infty}\left\lvert\frac{a_{n\nu}}{w_{n\nu-1}w_{n\nu-2}\cdots w_{0}}\right\rvert^{q}<\infty.$ Similarly, equating the coefficient of $k_{j},j=1,2,\cdots,\nu-1$ both sides we get $\sum_{n=1}^{\infty}\left\lvert\frac{a_{n\nu+j}}{w_{n\nu+j-1}w_{n\nu+j-2}\cdots w_{j}}\right\rvert^{q}<\infty.$ Consequently, the series $\sum_{n=1}^{\infty}\left\lvert\frac{a_{n}}{w_{0}w_{1}\cdots w_{n-1}}\right\rvert^{q}$ is convergent. It now remains to show that $\sum_{n=0}^{\infty}\lvert w_{-1}w_{-2}\cdots w_{-n}b_{-n-1}\rvert^{q}$ is convergent. To see that, we just need to equate the coefficient of $k_{-j-n\nu},j=0,1,\cdots,\nu-1$ both sides, and use the assumption $\sum_{n\geq 0}\left(\sum_{j\geq 1}\left\lvert\frac{\prod_{k=1}^{j}a_{-n-k+1}}{\prod_{k=1}^{j}b_{-n-k}}\right\rvert^{p}\right)^{\frac{q}{p}}<\infty$.
To get the implication in (iii) $\Rightarrow$ (i), we use the chaoticity criterion. As in the previous theorem, consider $D:=\text{span}\{k_{n}:n\in\mathbb{Z}\}$, and $S:D\rightarrow D$ which is given by $S(k_{n})=\frac{1}{w_{n}}k_{n+1},~{}~{}\forall~{}~{}n\in\mathbb{Z}.$ Observing that $F_{w}^{*}S(f)=f,$ $\forall$ $f\in D,$ we need only to prove that $\sum_{n=1}^{\infty}w_{-1}w_{-2}\cdots w_{-n}k_{-n}$ and $\sum_{n=1}^{\infty}\frac{1}{w_{0}w_{1}\cdots w_{n-1}}k_{n}$ are unconditionally convergent in $(\ell^{p}_{a,b}(\mathtt{A}))^{*}.$ To this end, since $k_{n}=a_{n}f_{n}^{*}+b_{n-1}f_{n-1}^{*}$, for all $n\in\mathbb{Z}$, we have
@@
As $\limsup_{n}|b_{n}/a_{n+1}|<1$, one gets $N_{1}\in\mathbb{N}$ and $r<1$ such that $|b_{n}/a_{n+1}|<r$ for all $n\geq N_{1}$. Hence,
@@
for all $n\geq N_{1}$. By our hypothesis, the right hand side of the above inequality is $q$-absolutely summable, and so, the series $\sum_{n\geq 1}\frac{1}{w_{0}w_{1}\cdots w_{n-1}}k_{n}$ is convergent. The unconditional convergence follows because $\{f_{n}^{*}\}$ is an unconditional basis. Similarly, to show that $\sum_{n=1}^{\infty}w_{-1}w_{-2}\cdots w_{-n}k_{-n}$ is unconditionally convergent, we write
@@
As $\limsup_{n}|a_{-n}/b_{-n-1}|<1$, we find $N_{2}\in\mathbb{N}$ and $R<1$ such that $|a_{-n}/b_{-n-1}|<R$ for all $n\geq N_{2}$. Hence,
@@
for all $n\geq N_{2}$. We can conclude that series $\sum_{n\geq 1}\frac{1}{w_{-1}w_{-2}\cdots w_{-n}}k_{-n}$ is unconditionally convergent. Hence, $F_{w}^{*}$ satisfies the chaoticity criterion, and (i) follows. ∎
The proofs of the main theorems on the hypercyclicity, supercyclicity, and chaos of $F_{w}^{*}$ on the dual of $\ell^{p}_{a,b}(\mathtt{A})$ show that we can have general results, stated below.
Theorem 3.13.
Let $\mathcal{E}$ be a Banach space of analytic functions on an annulus $\mathtt{A}$ concentric at 00, having bounded evaluation functionals. Let $\mathcal{E}^{*}$ be norm-separable, and let $\{k_{n}\}$ be linearly independent. If $F_{w}$ is bounded on $\mathcal{E}$, then the following hold.
  1. (i)
    $F_{w}^{*}$ is hypercyclic on $\mathcal{E}^{*}$ if
    $\liminf_{n\rightarrow\infty}~{}\Big{(}\big{\lvert}\prod_{j=1}^{n}w_{\nu-j}\big{\rvert}\|k_{\nu-n}\|\Big{)}=0=\liminf_{n\rightarrow\infty}~{}\Big{(}\frac{\|k_{\nu+n}\|}{\Big{\lvert}\prod_{j=0}^{n-1}w_{\nu+j}\Big{\rvert}}\Big{)},~{}~{}\forall~{}~{}\nu\in\mathbb{N}.$ (3.8)
  2. (ii)
    $F_{w}^{*}$ is mixing if
    $\lim_{n\rightarrow\infty}~{}\Big{(}\big{\lvert}\prod_{j=1}^{n}w_{\nu-j}\big{\rvert}\|k_{\nu-n}\|\Big{)}=0=\lim_{n\rightarrow\infty}~{}\Big{(}\frac{\|k_{\nu+n}\|}{\left\lvert\prod_{j=0}^{n-1}w_{\nu+j}\right\rvert}\Big{)},~{}~{}\forall~{}~{}\nu\in\mathbb{N}.$ (3.9)
  3. (iii)
    $F_{w}^{*}$ is supercyclic if
    $\liminf_{n\rightarrow\infty}\Big{(}\left|\frac{\prod_{j=1}^{n}w_{\nu-j}}{\prod_{j=0}^{n-1}w_{\nu+j}}\right|\|k_{\nu-n}\|\|k_{\nu+n}\|\Big{)}=0,~{}~{}\forall~{}~{}\nu\in\mathbb{N}.$
  4. (iv)
    Denote by $\widehat{f}(n)$, the $n$-th Laurent coefficient of $f\in\mathcal{E}$. If $\mathcal{E}$ is reflexive, then $F_{w}^{*}$ is chaotic provided
    $\sum_{n\geq 1}\Big{|}\Big{(}\prod_{j=1}^{n}w_{-j}\Big{)}\widehat{f}(-n)\Big{|}<\infty\hskip 11.38092pt\text{and}\hskip 11.38092pt\sum_{n\geq 1}\Big{|}\Big{(}\frac{1}{\prod_{j=1}^{n}w_{j}}\Big{)}\widehat{f}(n+1)\Big{|}<\infty,~{}~{}\forall~{}f\in\mathcal{E}.$
Proof.
Only an outline of the proofs are given. For all the implications, let $D$ and $S$ be as in the previous theorem. By a standard procedure, the results in (i), (ii), and (iii) follow from the hypercyclicity and supercyclicity criteria.
To see that (iv) also holds, via the chaoticity criterion we show that $\sum_{n=1}^{\infty}w_{-1}w_{-2}\cdots w_{-n}k_{-n}$ and $\sum_{n=1}^{\infty}\frac{1}{w_{1}w_{1}\cdots w_{n}}k_{n+1}$ are norm unconditionally convergent in $\mathcal{E}^{*}$. By the well known Orlicz-Pettis theorem (cf. [16]) it suffices to prove that these series are weakly unconditionally convergent. This is exactly the summability assumption in (iv). ∎
Next, we provide some necessary conditions for hypercyclicity, supercyclicity, and mixing of $F_{w}$ and its adjoint. We require the following basic result about the orbits under adjoints of hypercyclic and mixing operators.
Proposition 3.14.
(Bonet [6]) Let $T$ be a bounded operator on a Banach space $X$, and $x^{*}\neq 0$ be a bounded functional. If $T$ is hypercyclic, then $\sup_{n}\|T^{*n}x^{*}\|=\infty$.
If $T$ is mixing, then $\lim_{n\rightarrow\infty}\|T^{*n}x^{*}\|=\infty$.
We have:
Proposition 3.15.
Suppose that $F_{w}$ is bounded on a Banach space $\mathcal{E}$ of analytic functions over an annulus, having bounded evaluation functionals. For each fixed $\nu\in\mathbb{Z}$, the following hold.
  1. (i)
    If $F_{w}$ is hypercyclic, then $\sup_{n\geq 1}~{}(|w_{\nu-1}\cdots w_{\nu-n}|\|k_{\nu-n}\|)=\infty$. Also, if $F_{w}$ is mixing, then $\lim_{n\rightarrow\infty}~{}(|w_{\nu-1}\cdots w_{\nu-n}|\|k_{\nu-n}\|)=\infty.$
  2. (ii)
    Assume that $\mathcal{E}$ contains all Laurent polynomials. If the adjoint $F_{w}^{*}$ is hypercyclic,
    then $\sup_{n\geq 1}(|w_{\nu}\cdots w_{\nu+n}|\|z^{n+\nu}\|)=\infty.$ If $F_{w}^{*}$ is a mixing operator,
    then $\lim_{n\rightarrow\infty}(|w_{\nu}\cdots w_{\nu+n}|\|z^{n+\nu}\|)=\infty$.
Proof.
If $F_{w}$ is hypercyclic, then by the aforementioned result, we have $\sup_{n\geq 0}\|F_{w}^{*n}(k_{\nu})\|=\infty$. In view of Proposition 3.1, we have $F_{w}^{*n}(k_{\nu})=w_{\nu-1}\cdots w_{\nu-n}k_{\nu-n}$. So, the first part in (i) follows. Similarly, the other part also holds.
If $F_{w}^{*}$ is hypercyclic, then the orbit $\{F_{w}^{n}(z^{\nu}):n\geq 0\}$ is unbounded. Noting that $F_{w}^{n}(z^{\nu})=w_{\nu}\cdots w_{\nu+n}z^{\nu+n}$, we get (ii). ∎
Next, we state a result on the existence of hypercyclic subspaces for $F_{w}$ and $F_{w}^{*}$. It is well known that an operator $T$ on a complex Banach space, satisfying the hypercyclicity criterion has a hypercyclic subspace if and only if
$\sigma_{e}(T)\cap S^{1}\neq\phi,$
$S^{1}$ being the unit circle in $\mathbb{C}$. Every hypercyclic weighted shift on $\ell^{p}(\mathbb{Z})$ has a hypercyclic subspace. This is not true for hypercyclic unilateral shifts, and a characterization along with other results can be found in [4], [27], [29], and [30].
Theorem 3.16.
Suppose $F_{w}$ satisfies the conditions of Theorem 3.6. If $F_{w}$ is hypercyclic, then it has hypercyclic subspaces. A similar conclusion holds for $F_{w}^{*}$.
@@ 3.3. Examples and Illustrations
We provide explicit examples to illustrate our results on the linear dynamics of $F_{w}$ and $F_{w}^{*}$. Recall that a Laurent series $\sum_{n=-\infty}^{\infty}c_{n}z^{n}$ converges if $r<\lvert z\rvert<R,$ where
@@
Indeed, $|z|<R$ and $\frac{1}{|z|}<\frac{1}{r}$ give, respectively, the convergence of $\sum_{n\geq 0}c_{n}z^{n}$ and $\sum_{n\leq 0}c_{n}z^{n}$.
@@
Let $a_{0}=b_{0}=w_{0}=1,$ and $a_{n}=1,$ $b_{n}=\frac{1}{n},$ $w_{n}=\frac{1}{(n+1)!},$ for all $n\geq 1$, along with $a_{n}=n2^{n-2},$ $b_{n}=n2^{n},$ $w_{n}=8,$ for all $n\leq-1.$ By Theorem 2.1, for $1<p<\infty,$
$\sum_{n=-\infty}^{\infty}\left((\lvert a_{n}\rvert+\lvert b_{n}\rvert)\lvert z\rvert^{n}\right)^{q}=2^{q}+\sum_{n=1}^{\infty}\left(\left(\frac{n}{2^{n}}\frac{1}{2^{2}}+\frac{n}{2^{n}}\right)\lvert z\rvert^{-n}\right)^{q}+\sum_{n=1}^{\infty}\left(\left(1+\frac{1}{n}\right)\lvert z\rvert^{n}\right)^{q},$
where $11/p+1/q=$ 1. A simple computation shows that the natural annulus $\mathtt{A}$ of $\ell^{p}_{a,b}(\mathtt{A})$ is $r<|z|<R$, where $r=\frac{1}{2}$ and $R=1.$ Therefore $\sum_{n=-\infty}^{\infty}\left((\lvert a_{n}\rvert+\lvert b_{n}\rvert)\lvert z\rvert^{n}\right)^{q}<\infty$ if $\frac{1}{2}<\lvert z\rvert<1.$ In view of Theorem 3.3, $F_{w}$ is bounded on $\ell^{p}_{a,b}(\mathtt{A}),$ where $\mathtt{A}$ is the given by annulus $\frac{1}{2}<\lvert z\rvert<1.$ Additionally, $F_{w}$ satisfies the conditions of Theorems 3.7 and 3.9, and so $F_{w}$ is mixing and chaotic on $\ell^{p}_{a,b}(\mathtt{A}).$ Indeed,
$\lim_{n\rightarrow\infty}\left\lvert\prod_{k=0}^{n-1}w_{k}^{-1}\prod_{k=0}^{n-1}\frac{b_{k}}{a_{k}}\right\rvert=\lim_{n\rightarrow\infty}\frac{\prod_{k=1}^{n}k!}{(n-1)!}=\infty~{}~{}~{}~{}\text{and}~{}~{}~{}~{}\lim_{n\rightarrow\infty}\left\lvert\prod_{k=1}^{n}w_{-k}\prod_{k=1}^{n}\frac{a_{-k}}{b_{-k}}\right\rvert=\lim_{n\to\infty}2^{n}=\infty.$
@@
In the previous example, if we let $w_{n}=1,$ for all $n\leq-1,$ $w_{n}=\frac{1}{(n+1)!},$ for all $n\geq 0,$ and $\{a_{n}\},\{b_{n}\}$ being the same as mentioned above, then we see that by Theorem 3.7, $F_{w}$ is supercyclic. Now, by Theorem 3.15, we know that if $F_{w}$ is hypercyclic on $\ell^{p}_{a,b}(\mathtt{A}),$ then $\sup_{n\geq 1}~{}(|w_{-1}\cdots w_{-n}|\|k_{-n}\|)=\infty$, which suggests that $\sup_{n\geq 1}~{}(|w_{-1}\cdots w_{-n}b_{-n-1}|)=\infty$, by using Proposition 2.2 and $\lim\sup_{n\rightarrow+\infty}\big{|}\frac{a_{-n+1}}{b_{-n}}\big{|}<1.$ However, in this case $F_{w}$ is not hypercyclic on $\ell^{p}_{a,b}(\mathtt{A})$ as
$\sup_{n\geq 1}\left\lvert b_{-n-1}\prod_{k=1}^{n}w_{-k}\right\rvert=\sup_{n\geq 0}\frac{n+1}{2^{n+1}}<\infty.$
@@
$\sum_{n=-\infty}^{\infty}\left((\lvert a_{n}\rvert+\lvert b_{n}\rvert)\lvert z\rvert^{n}\right)^{q}=2^{q}+\sum_{n=1}^{\infty}\left(\left(\frac{1}{n}+1\right)\lvert z\rvert^{-n}\right)^{q}+\sum_{n=1}^{\infty}\left(\left(\frac{n}{2^{n}}+\frac{n}{2^{n+2}}\right)\lvert z\rvert^{n}\right)^{q},$
where $11/p+1/q=$ 1. In this case, $r=1$, $R=2$, and hence, the annulus $\mathtt{A}$ is $21<|z|<$ 2. Further, $\sum_{n=-\infty}^{\infty}\left((\lvert a_{n}\rvert+\lvert b_{n}\rvert)\lvert z\rvert^{n}\right)^{q}<\infty$ if $21<\lvert z\rvert<$ 2 . In addition, by the conditions in Theorem 3.3, the weighted shift $F_{w}$ is bounded on $\ell^{p}_{a,b}(\mathtt{A}).$ In view of Theorems 3.10 and 3.12, the adjoint $F_{w}^{*}$ is mixing and chaotic on $(\ell^{p}_{a,b}(\mathtt{A}))^{*}.$ Indeed,
$\lim_{n\rightarrow\infty}\left\lvert b_{-n-1}\prod_{k=1}^{n}w_{-k}\right\rvert=\lim_{n\rightarrow\infty}\frac{1}{\prod_{k=3}^{n+2}k!}=0~{}~{}~{}~{}\text{and}~{}~{}~{}~{}\lim_{n\rightarrow\infty}\left\lvert~{}\frac{a_{n}}{\prod_{k=0}^{n-1}w_{k}}\right\rvert=\lim_{n\rightarrow\infty}\frac{n}{2^{4n-3}}=0.$
Remark 3.20.
It should be noted that the operators $F_{w}$ considered in this section satisfy
$\lim_{n\rightarrow\infty}\left\lvert w_{n+1}\frac{b_{n}}{a_{n+2}}-w_{n}\frac{a_{n}}{a_{n+1}}\frac{b_{n+1}}{a_{n+2}}\right\rvert=0~{}\text{and}~{}\lim_{n\rightarrow-\infty}\left\lvert w_{n}\frac{a_{n}}{b_{n}}-w_{n+1}\frac{a_{n+1}}{b_{n+1}}\right\rvert=0.$
Therefore, by the Theorem 3.6, we see that $F_{w}$ on $\ell^{p}_{a,b}(\mathtt{A})$ is similar to a compact perturbation of a bilateral weighted forward shift on $\ell^{p}(\mathbb{Z}).$ Consequently, we have a class of chaotic operators that are compact perturbation of some classical weighted shifts.